a)     Define marginal utility. Provide an argument why to maximize total utility of good x and good y, the consumer should consume until the ratio of marginal utilities over price is the same across both goods. 

                                                                      


1
Expert's answer
2021-05-06T16:08:15-0400

Marginal utility is the additional satisfaction derived from having one more unit of a good or service.

The consumer equilibrium is the situation when consumer maximizes satisfaction by consuming the goods by spending all the income. It is a situation when marginal utility of good and price of the good is equal for all the commodities. In case of two goods, it is indicated as follows:

"\frac{MU_x}{P_x }= \frac{MU_y}{P_y}"

As per law of diminishing marginal utility theory, the increment in the consumption of good on continuous basis reduces the marginal utility and raises the total utility at decreasing rate. 

If the marginal utility over price ratios are not equal then the satisfaction of the consumers is not maximized. 

Case-1: 

"\frac{MU_x}{P_x }>\frac{MU_y}{P_y}"

In this case, the consumption of good X must be increased and consumption of good y must decreased to make these ratios equal.

Case-2:

"\frac{MU_x}{P_x }< \frac{MU_y}{P_y}"

In this case, the consumption of good Y must be increased and consumption of good X must decreased to make these ratios equal.